# Examining Inequality in Park Quality for Promoting Health Across 35 Global Cities

### v1

@@ 4.1 Data
OpenStreetMap (OSM) is a globally encompassing geographic information database based on crowdsourced contributions. While accessible through a map interface at https://openstreetmap.org, its primary value lies in its data being an indispensable source for third-party mapping, navigation applications[182, 183], and diverse scientific analyses, such as urban planning[184, 185], geospatial analyses[186, 187], and disaster management[188, 189]. With its permissive licensing, the project has nurtured a vast ecosystem of contributors, from individuals to corporations engaged in the geographic information sector. As a result, OSM has achieved comprehensive worldwide coverage, with near-perfect mapping quality across the western world[190], while retaining remarkable detail in the global south[188].
For this study, we utilized an OSM data dump obtained from https://geofabrik.de in fall 2023. OSM uses three data structures to describe map objects — nodes, ways, and relations. Furthermore, OSM employs a tagging system with key-value pairs to categorize and describe all these map objects. Each map object is typically associated with multiple tags, i.e., key-value pairs, that provide insights into its main purpose, used primarily for rendering the object on the map. However, these tags can encompass various additional information, such as opening hours, inscriptions, and data source references.
To avoid the complexities of the OSM data model[191], for our tasks, it was sufficient to focus on two key map objects related to parks: park elements and park spaces. Park elements are 0-dimensional points representing objects like benches, individual trees, and statues. These elements are stored as “nodes” in the OSM data model. On the other hand, park spaces refer to areas within the parks, such as meadows, lakes, and forests. In the OSM database, these areas can be modeled as “ways” or “relations”, but for the purpose of this study, we treat all areas equally.
@@
We utilized a substantial dataset comprising geo-located images posted between 2004 and 2015. This extensive dataset offered us a valuable secondary perspective on activities taking place within the parks of the world. By intersecting these images with the park outlines from OSM, we identified 10,788,686 pictures captured within the boundaries of parks in our study cities. To extract the depicted content from these images, we used user-assigned tags in conjunction with automatically-generated computer vision labels [192, 193]. These labels enabled us to gather information about the various facilities and activities within the parks.
@@ 4.2 Study Area
Our research focuses on 35 cities listed in Table 6 selected using three criteria to make our analysis broad yet robust.
First, we chose to look at big cities worldwide with at least 650,000 people. This size was selected because it includes many of the largest cities, including most of the large capitals in Europe and other densely populated areas across the globe where parks are crucial for people’s well-being [9, 194]. It also helped keep our study focused and understandable without including too many smaller cities. However, to ensure better representation in Oceania, we made an exception for Christchurch in New Zealand, which we included, although it is less populous than our cutoff. This first criterion allowed us to examine parks in cities from various parts of the world, each affected by its own climate, history, and cultural background.
Second, we only looked at cities in countries where at least 80% of the population has access to the Internet [195]. This ensured we had enough online data (like tags on OSM or photos on Flickr) for our study. Since there is no detailed global data on Internet use in cities specifically, we used the country’s overall access to the Internet as our guide. This way, we set a basic standard, knowing that people in cities, especially in developing countries, usually have better access to the Internet than those in rural areas. We decided on this threshold upon our preliminary empirical analyses, finding that in many cities in Africa and South America, where fewer people have Internet, there was not enough digital information for our data-intensive approach.
Third, we chose cities where, on average, parks have at least one-eighth of their area tagged with health-related tags on OSM. Because we relied heavily on OSM for our analyses, this rule helped us focus on cities with sufficient information on the platform for our study. We settled on this one-eighth rule after observing that this was the threshold where the lack of manual tagging prevented us from obtaining sufficient information for our task. Mainly, this meant we could not include cities where the predominantly automatic tagging of areas in OSM through earth observation was not accompanied by manual tagging of OSM contributors. This was mainly the case in China, where OSM mapping is deemed illegal[196].
@@ 4.3 Identifying Health-promoting Activities in Urban Greenery
We used an expert panel to identify and categorize activities and discovered their health benefits through a scoping literature review.
Expert Panel
First, we assembled a list of activities that people engage in at urban parks. This was accomplished with the help of an expert panel composed of three co-authors of this paper. To collect relevant papers, we used two specific queries of Google Scholar: “(urban) AND (parks OR greenery) AND usage” and “(activities in urban) AND (parks OR greenery)”. From this process, we collected the top 50 scholarly articles for each search phrase. This resulted in 91 unique papers.
We carefully examined each article, making a note of all the activities mentioned. This gave us a wide range of activities that varied in their level of detail. For instance, both broad activity terms like leisure activities or recreation as well as more specific activity categories like physical and social activities were frequently mentioned. We also noted down plenty of individual activities like walking, performing street theatre, fishing, and playing all kinds of different sports. There were instances where we inferred activities from the mention of specific facilities. For example, the mention of playgrounds was noted as “playing”, boat rentals as “boating”, and fitness facilities as “exercising”. Afterward, we held a meeting to categorize these activities. We especially focused on the potential health benefits of each activity and tried to find the similar levels of detail for each category. For instance, leisure activities or recreation were seen as too broad categories, while distinguishing between physical and social activities was considered appropriate. In the end, six categories emerged: physical, mind-body, nature-appreciation, environmental, social, and cultural activities (Section 2.1, Step 1: Identifying Activities in Parks.).
Scoping Review
@@ 4.4 Annotating Park OSM Tags with Activities Using LLM Classifiers
To associate different park elements and spaces with health-promoting activities, we annotated OSM tags describing those elements and spaces with activities. This turned out to be a challenging task. OSM is a collaborative platform with some governance and guidelines[178] for tagging, but the flexible tagging system offers the crowdsourcing contributors substantial freedom in their tasks. Consequently, the data can be disorganized, requiring intense preprocessing and tidying up. Each map object, such as park elements and spaces, can be marked with an unlimited number of tags, offering in-depth descriptions. This results in a large variety of tags — specifically, we noted more than 30,000 unique key-value pairs describing park elements and spaces. Given that our main interest is in the primary aspect of the park elements and spaces, we performed a data cleaning step, detailed in the Appendix, Section B.2, where we eliminated irrelevant meta information of the items. This step lets us concentrate exclusively on the key aspects linked to health-promoting activities, making the labeling task more manageable and pertinent to our study.
@@
Even for domain experts, it was challenging to unequivocally link these tags to health-related activities. For example, a bench could be associated with various health benefits — socializing, enjoying nature, or resting from physical activity. Choosing one activity over another depended largely on personal experiences [198], as many tags could be linked to multiple activities. Given the vast number of items and the specialized nature of the task, we opted for an LLM classifier as an alternative to expert annotation or crowdsourcing.
Using LLMs as classifiers comes with several benefits. They offer a more objective and consistent way to annotate that can be applied to a large number of necessary annotations in a short time and at low costs. In some instances, these classifiers have even been found to perform better than crowdworkers[199]. This is particularly true as crowdworkers might themselves use available machine learning models for performing tasks[200]. Furthermore, research suggests that the quality of annotation achieved with LLM classifiers could be on par with domain experts [201]. Therefore, we chose to establish a benchmark in our domain to evaluate the feasibility of using LLM annotations in mapping OSM tags to health-promoting activities. To do so, we had three of our experts manually annotate a set of 100 most frequent tags, and then we chose the final labels with majority voting. We then used this expert-labeled dataset to assess the accuracy of the labels generated by different LLM classifiers. The details of the LLM annotation benchmark can be found in the Appendix, Section B.3. The outcome indicated that GPT-4, set at a temperature of $0.9$, yielded the best annotation performance of an F1 score of $0.77$.
@@
Using the taxonomy with six categories of health-promoting activities in Table 1, and GPT-4 as the best-performing annotation model, we ran the annotation of OSM tags describing park elements and park spaces. These tags were then labeled with one of the health-promoting activities, or “none” if they didn’t support a particular activity. By doing so, we gained insights into which type of health-promoting activities each park element and park space potentially facilitate, which in turn led to a comprehensive and precise understanding of the health-promoting activities in parks and their potential contributions to public health (Table 2). Note that the mind-body activities category was discarded in this step since none of the OSM tags found in parks were primarily mapped to this category.
@@ 4.5 Computing Park Health Scores by Aggregating OSM Tags
In our process of assigning health-promoting activity scores to each park, we first gathered park elements and spaces within each park. We used the osmium library to collect all relevant map resources from OSM that intersected with the shape of each park. For park elements, we simply counted all elements found within the park boundaries. But for the spaces, we calculated the overlapping area with the park’s area. This was important because some spaces, like a river running through the park, might continue beyond the park’s edges and should not fully count towards a park’s score. After obtaining two lists – one for park elements and another for park spaces – we then assigned health-promoting activities to these elements and spaces based on the LLM classification of their associated tags, such as shown in Table 2. We discarded any elements or spaces whose tags did not match an activity category. In a few instances, park elements or spaces might have fit into more than one health-promoting activity category. For example, sports fields were often tagged with different labels that contribute to both physical and social categories. In these cases, we proportionally counted the resource, taking into account all the matched categories under which it falls.
@@
After tallying up the park elements and spaces within the park, we measured the overall effect of the park in promoting healthy activities within a city. This score should account for the park’s size and the range of facilities it offers for different activities. Our proposed scoring method is based on these considerations.
@@
    Density of Health-Promoting Elements and Spaces: The character of a park is determined by the concentration of health-promoting facilities.
@@
    City-specific Normalization of Park Health Scores: The value of a single park’s facilities for a certain activity is relative to similar facilities in other parks throughout the city.
@@
To reflect these assumptions into our scoring, we proposed a linear regression model to compute the park health scores shown in Equation 1. We used the idea of an “average park” in each city to compute a baseline and used the distance of each park to the average park line, i.e., the residual, as score. The average park baseline was determined by computing separate linear regression models for park elements and spaces in each city, estimating the expected amount of facilities relative to the park area.
@@
$E_{Act}(log_{2}(count(Act)))=i+s\cdot log_{2}(\text{park area})\leavevmode\nobreak\ |\leavevmode\nobreak\ Act\in\text{Activity Categories},$ (1)
where $i$ and $s$ represent the intercept and slope of the regression lines, respectively. To obtain regression models for each activity category and both park elements and park spaces, we utilized the binary logarithm to account for the diminishing returns of an increase in park size. For each city, separate regression models were calculated for each activity category, as well as for park elements and park spaces. During the computation of the regression models, parks with very low activity counts in a specific category were excluded. This exclusion was necessary to prevent artificially flattening the regression lines due to close-to-zero values, which would distort the normalization. The specific threshold for excluding parks with low activity counts was determined empirically by analyzing the histograms of the values. This approach enabled us to identify an appropriate cutoff point for excluding parks with insufficient activity data, ensuring the reliability of the regression models. For a visual representation of the exclusion process and the determination of the threshold, refer to Figure 12 and Figure 13 in the Appendix.
By considering the residuals, we could identify parks that exceeded expectations (positive residuals) and parks that fell short in terms of providing health-promoting resources (negative residuals) for this activity. Making the linear model dependent on the park area, we achieved that the resulting scores of health-promoting elements and spaces depend on the density. We used the binary logarithm to discount the effects of extremely large counts and parks. Furthermore, we chose to calculate these scores for each city separately rather than using a single global model. This way, we ensured that the benefits accurately represented each city’s local situation. To visualize this method, we plotted the linear models and the individual park scores in the $log-log$ space for London, UK, in Figure 7. The regression lines denote expected health scores based on park size. Park scores for park elements and spaces are residuals from this average park line in the model, adjusting for park size when determining health scores.
@@
Figure 7: Visualization of the scoring method. The linear regression models for determining the park health scores for the park elements (top) and park spaces (bottom) in London, UK. The horizontal axis denotes the of the park’s area ($log_{2}$), while the vertical axis represents of the count of park elements and the area occupied by health-promoting park spaces ($log_{2}$). The modest $R^{2}$ values are anticipated, highlighting the variety among parks.
@@
The regression models gave us individual scores for park elements and park spaces for each activity. To unify these scores into one combined health score, we examined the co-occurrences of park elements and spaces and found that they represent orthogonal concepts in practice. Park elements include points of interest, individual trees, benches, and similar items. Park spaces, however, include areas like forests, sports fields, and buildings. There can be cases where a park area is broken down into its individual parts, like a playground with separately mapped features like swings, slides, or spinning equipment. But these cases are pretty rare in OSM mapping. Likewise, unless a tree is particularly important, areas tagged as natural=wood should not include individual trees according to the mapping guidelines[178]. Based on these observations, we hypothesized that it would be acceptable to combine scores from park features and areas linearly, as they contribute differently to the overall offering of facilities for health-promoting activities.
Activity Category Mean Pearson Correlation
Physical 0.16
Nature-appreciation 0.05
Environmental 0.18
Social 0.06
Cultural 0.18
To achieve this, we first normalized the scores of park elements and spaces using the z-score transformation considering all parks in a city. This normalization allowed us to standardize the scores, making them comparable despite being on different scales. Then, we linearly combined these z-scores by averaging them together to create the overall score for the park according to Equation 2.
@@
where $P$ denotes an individual park, $Act$ is one of the activity categories, the residual scores for element and spaces stem from Equation 1, and $z()$ indicates the z-score transformation.
The combined and normalized scores of park elements and spaces represent a comprehensive and unified measure of the park’s health-promoting amenities and facilities, accounting for both individual elements and cultivated areas. The combination process accounts for the relative importance of each aspect, leading to a more meaningful overall score that represents the park’s overall quality in terms of offerings toward health-promoting activities.

### v2

@@ 4.1 Data
OSM is a globally encompassing geographic information database based on crowdsourced contributions. While accessible through a map interface at https://openstreetmap.org, its primary value lies in being an indispensable source for open mapping data both in commercial and scientific applications [54, 55, 56]. Thanks to its permissive licensing, OSM has fostered a large ecosystem of individual and professional contributors. As a result, the project has achieved comprehensive worldwide coverage, with near-perfect mapping quality across the western world[56], while retaining remarkable detail in the global south[55]. In this study, we utilized OSM data from fall 2023.
OSM employs a tagging system with key-value pairs to categorize and describe all these map objects. Each map object is typically associated with multiple tags that describe its purpose, but may also include additional information, such as opening hours, or data source references. To avoid the complexities of the OSM data model, for our tasks, it was sufficient to focus on two key map objects related to parks: park elements and park spaces. Park elements are 0-dimensional points representing objects like benches, individual trees, and statues. On the other hand, park spaces refer to areas within the parks, such as meadows, lakes, and forests.
@@
We utilized a substantial dataset comprising geo-located images posted between 2004 and 2015. This extensive dataset offered us a valuable secondary perspective on activities taking place within the parks of the world. By intersecting these images with the park outlines from OSM, we identified 10,788,686 pictures captured within the boundaries of parks in our study cities. To extract the depicted content from these images, we used user-assigned tags in conjunction with automatically-generated computer vision labels [57, 58].
@@ 4.2 Study Area
Our research focused on 35 cities listed in Table 1, which we selected using three criteria to make our analysis broad yet robust.
First, we selected major cities worldwide with populations of at least 650,000. This threshold includes many of the largest urban areas, such as major European capitals and other densely populated regions where parks play a vital role in public well-being. To improve representation in Oceania, however, we made an exception for Christchurch, New Zealand, which has a smaller population. This first criterion allowed us to examine parks in cities from various parts of the world, each affected by its own climate, history, and cultural background.
Second, we only looked at cities in countries where at least 80% of the population has access to the Internet [59]. This ensured we had enough online data (like tags on OSM or photos on Flickr) for our study. Since there is no detailed global data on Internet use in cities specifically, we used the country’s overall access to the Internet as our guide. We decided on this threshold upon our preliminary analyses, finding that in many cities in Africa and South America, there was not enough digital information for our approach, which relies on social media and collaborative mapping data.
Third, we chose cities where, on average, parks have at least one-eighth of park areas are annotated with health-related tags on OSM. Since our analysis relied heavily on OSM data, this criterion ensured a minimum level of information on the platform for our study. We settled on this one-eighth threshold after observing that, below this level, the lack of contributor-added tags limited our ability to extract meaningful information. This primarily excluded cities where most tagging was done predominantly automatically through earth observation that was not accompanied by manual tagging of OSM contributors. This was mainly the case in China, where non-governmental mapping is restricted[60].
@@ 4.3 Identifying Health-promoting Activities in Urban Greenery
We identified and categorized park-based activities using input from an expert panel consisting of three co-authors of this study. We compiled a comprehensive list of activities commonly undertaken in urban parks. To collect relevant papers, we used two specific queries of Google Scholar: “(urban) AND (parks OR greenery) AND usage” and “(activities in urban) AND (parks OR greenery)”. From this process, we retrieved the top 50 scholarly articles for each search phrase, resulting in a total of 91 unique papers. We reviewed each article and collected all activities, resulting in a diverse set of activity descriptions varying in granularity. For example, the literature included both broad terms such as leisure activities or recreation, as well as more specific categories like physical and social activities. We also noted plenty of individual activities like walking, performing street theater, fishing, and playing all kinds of different sports. Subsequently, we convened to categorize the identified activities, with a particular focus on their potentially different impact on health and on ensuring a consistent level of specificity across categories. Broad terms such as recreation were deemed too general to be analytically useful, whereas distinctions like physical versus social activities were considered meaningful. This process ultimately yielded six distinct categories: physical, mindfulness, nature-appreciation, environmental, social, and cultural activities.
@@ 4.4 Annotating Park OSM Tags with Activities Using LLM Classifiers
To associate different park elements and spaces with health-promoting activities, we annotated OSM tags describing those elements and spaces with activities. This turned out to be a challenging task. OSM is a collaborative platform with some governance and guidelines (https://wiki.openstreetmap.org/wiki/Map_features) for tagging, but the flexible tagging system offers the crowdsourcing contributors substantial freedom. As a result, the data can be inconsistent and fragmented, necessitating thorough cleaning. Each map object, such as park elements and spaces, can be tagged with an unlimited number of tags, offering in-depth descriptions. As a result, we encountered over 30,000 unique key–value pairs associated with park elements and spaces. Since our primary focus is on the core functional aspects of these elements, we conducted a data cleaning step (detailed in Appendix Section B.3) to remove irrelevant metadata associated with the map objects. This filtering allowed us to focus exclusively on tags relevant to activity-related features, thereby making the annotation process more pertinent to our study.
@@
Even for domain experts, linking these tags unequivocally to health-related activities was difficult. For instance, a bench might relate to socializing, enjoying nature, or resting after physical activity. Choosing one activity over another often depended on personal experience, as many tags could plausibly refer to multiple activities. Given the large number of items and the specialized nature of the task, we chose an LLM classifier as an alternative to expert annotation or crowdsourcing.
Using LLMs as classifiers offers several advantages as they provide a more objective and consistent approach to annotation, can handle large volumes of data quickly, and do so at relatively low cost. In some cases, LLM classifiers have even outperformed crowdworkers[61], who may themselves rely on machine learning tools to complete tasks[62]. Research also suggests that LLM-based annotations can match the quality of those made by domain experts[63]. Given these benefits, we established a benchmark in our domain to evaluate the feasibility of using LLMs to map OSM tags to health-promoting activities. Three experts manually annotated the 100 most frequent tags, and final labels were selected by majority vote. We then used this expert-labeled dataset to assess the accuracy of the labels generated by different LLM classifiers. The details of the LLM annotation benchmark can be found in the Appendix, Section B.4. The outcome indicated that GPT-4, set at a temperature of $0.9$, yielded the best annotation performance of an F1 score of $0.77$.
@@
Using the taxonomy with six categories of health-promoting activities in Table 7, and GPT-4 as the best-performing annotation model, we ran the annotation of OSM tags describing park elements and park spaces. These tags were then labeled with one of the health-promoting activities, or “none” if they didn’t support a particular activity. In doing so, we established a lexicon of park elements and spaces linked to health-promoting activities (Table 3). We had to exclude the mindfulness activities category at this stage, as none of the OSM tags found in parks were primarily associated with it.
@@ 4.5 Computing Park Health Scores by Aggregating OSM Tags
In our process of assigning health-promoting activity scores to each park, we first gathered park elements and spaces within each park using the osmium library. We then assigned health-promoting activities to these elements and spaces based on the lexicon created in the previous step (Table 3). We discarded any elements or spaces whose tags did not match an activity category. In a few instances, park elements or spaces could fall into more than one health-promoting activity category. For example, apple trees are annotated in OSM with the tags [natural=tree, produce=apple]. In our lexicon of park elements and spaces, we map natural=tree to the nature-appreciation and produce=apple to the environmental category. To account for this overlap, we proportionally assign the element as 50% nature-appreciation and 50% environmental. More generally, when multiple tags are matched to different categories, we proportionally count the resource based on the number of matched tags, ensuring that its contributions are accurately accounted for, thus avoiding underestimating secondary activities of multi-purpose facilities.
@@
After tallying up the park elements and spaces within the park, we measured the overall effect of the park in promoting healthy activities within a city. This score should account for the park’s size and the range of facilities it offers for different activities. Our proposed scoring method is based on the following considerations.
@@
    Amount of Health-Promoting Elements and Spaces: The number and area of health-promoting facilities determine how much offering there is for each activity.
@@
    Area of the Park: The character of a park depends on the concentration of health-promoting facilities. Larger parks must offer more to obtain a high score.
  3. 3.
@@
    City-specific Normalization of Park Health Scores: The value of a park’s facilities for a certain activity is relative to similar facilities in other parks throughout the city.
@@
To reflect these assumptions into our scoring, we proposed a linear regression model to compute the park health scores shown in Equation 1. We used the idea of an “average park” in each city to compute a baseline and used the distance of each park to the average park line, i.e., the residual, as a score. The average park baseline was determined by computing separate linear regression models for park elements and spaces in each city, estimating the expected amount of facilities relative to the park area.
@@
$E_{Act}(log_{2}(count(Act)))=i+s\cdot log_{2}(\text{park area})\penalty 10000\ |\penalty 10000\ Act\in\text{Activity Categories},$ (1)
where $i$ and $s$ represent the intercept and slope of the regression lines, respectively. To obtain regression models for each activity category and both park elements and park spaces, we utilized the binary logarithm to account for the diminishing returns of an increase in park size. For each city, separate regression models were calculated for each activity category, as well as for park elements and park spaces. See Section B.5 for details on the modeling.
@@
Figure 4: Visualization of the scoring method. The linear regression models for determining the park health scores for the park elements (top) and park spaces (bottom) in London, UK. The horizontal axis denotes the park’s area ($log_{2}$), while the vertical axis represents the count of park elements ($log_{2}$) and the area occupied by health-promoting park spaces ($log_{2}$). The modest $R^{2}$ values are anticipated, highlighting the variety among parks.
By analyzing the residuals, we identified parks that exceeded expectations (positive residuals) and those that fell short (negative residuals) in providing health-promoting resources for a given activity. We made the linear model dependent on park area so that the resulting scores for health-promoting elements and spaces reflected their density. To reduce the influence of extremely large parks with high amounts of facilities, we applied the binary logarithm. We calculated these scores separately for each city, rather than using a single global model, to ensure that the results reflected each city’s local context. To illustrate this method, we plotted the linear models and the individual park scores in the loglogloglog space for London, UK (Figure 4). The regression lines denote expected health scores based on park size. The modest $R^{2}$ values showed that the number of health-promoting facilities could not be explained by park size but instead reflected different design priorities and the needs of local citizens. Park scores for park elements and spaces are residuals from this average park line in the model, adjusting for park size when determining health scores.
@@
The regression models gave us individual scores for park elements and park spaces for each activity. To verify the impact of combining these scores into one combined health score, we examined the co-occurrences of park elements and spaces and found that they represent orthogonal concepts in practice, as outlined in Section B.6. Based on this finding, we computed a combined score by first normalizing the scores of park elements and spaces using the z-score transformation, considering all parks in a city. This normalization allowed us to standardize the scores, making them comparable despite being on different scales. Then, we linearly combined these $z$-scores by averaging them together to create the overall score for the park according to Equation 2.
@@
where $P$ denotes an individual park, $Act$ is one of the activity categories, the residual scores for elements and spaces stem from Equation 1, and $z()$ indicates the $z$-score transformation.
The combined and normalized scores of park elements and spaces represent a comprehensive and unified measure of the park’s health-promoting amenities and facilities, accounting for both individual elements and cultivated areas. The combination process accounted for the relative importance of each aspect, leading to a more meaningful overall score that represents how well a park is equipped to support performing health-promoting activities. Since the scores are based on $z$-score–normalized residuals, a value around 0 indicates average support for a given activity, while a score of $\pm 1$ means the park is $1$ standard deviation above or below the city-wide average.